OpenStudy (anonymous):

What is the solution set of -9x ≥ 27?

OpenStudy (ganpat):

-9x ≥ 27 can u isolate x ??

OpenStudy (ganpat):

i mean, keep x on one side and rest all on other.. can u do that ??

OpenStudy (ganpat):

x ≥ 27 / (-9) So x ≥ -3.... is the solution...

OpenStudy (ganpat):

not getting ??

OpenStudy (maheshmeghwal9):

\[-9x \ge 27.\]\[\implies -x \ge \frac{27}{9}.\]\[\implies -x \ge 3.\]\[\implies x \le -3.\]Therefore; \[x \in (-\infty,-3].\] Becoz inequality reverses itself when multiplied or divided by a negative number on both sides of inequality.

OpenStudy (ganpat):

oh yes.. i missed negative sign... !!

OpenStudy (ganpat):

thanks, @maheshmeghwal9 ..

OpenStudy (maheshmeghwal9):

np:)